OpenStudy (lonewolf):

If p varies directly with q, and p = 8 when q = 2, what is the value of p when q = 7? A.p=40 B.p=28 C.p=21 D.p=100

sam (.sam.):

Varies directly means p=kq, \(\frac{p}{q}=k\)

sam (.sam.):

So \(k=\frac{8}{2}=4\) Now when q=7, and we've got k, p=kq p=(4)(7)=?

OpenStudy (lonewolf):

oh thank you

OpenStudy (lonewolf):

wow you are really smart, wish i was

sam (.sam.):

Yw

OpenStudy (lonewolf):

wait i have another question would this have the same formula as the one before?

OpenStudy (lonewolf):

The value of y varies directly with x2, and y = 64 when x = 4. What is the value of y when x = 6?

OpenStudy (lonewolf):

all i need to know if it still has the same formula